OpenStudy (anonymous):

determine the inverse of g(x) = x/1-x and the domain and range of g(x) and inverse of g(x)

OpenStudy (gorv):

domain all real number except x=1 range real numbers for inverse g(x)=x/(1-x) g(x)*(1-x)=x g(x)-x*g(x)=x g(x)=x+x*g(x) g(x)=x*(1+g(x)) g(x)/(1+g(x))=x replace x by g^-1(x) x/(1+x)=g(x) inverse